# A Fundamental Trade-off in Aligned Language Models and its Relation to Sampling Adaptors

### v1

@@ Abstract
The relationship between the quality of a string and its probability ${p}({\boldsymbol{y}})$ under a language model has been influential in the development of techniques to build good text generation systems. For example, several decoding algorithms have been motivated to manipulate ${p}({\boldsymbol{y}})$ to produce higher-quality text. In this work, we examine the probability–quality relationship in language models explicitly aligned to human preferences, e.g., through Reinforcement Learning through Human Feedback (RLHF). We find that, given a general language model and its aligned version, for corpora sampled from an aligned language model, there exists a trade-off between the average reward and average log-likelihood of the strings under the general language model. We provide a formal treatment of this issue and demonstrate how a choice of sampling adaptor allows for a selection of how much likelihood we exchange for the reward.
@@ 1 Introduction
The relationship between the probability of a string and its quality as judged by a human, is fundamental to our ability to deploy language models that generate useful text. The probability of a string ${\boldsymbol{y}}$ under a language model ${p}$ is commonly used as a heuristic to reason about text quality. The intuition behind this approach is that under a well-calibrated language model trained primarily on human-written text, strings that occur with high probability should be more human-like, and thus judged by humans to be of higher quality. In other words, a priori, one should expect a positive correlation between a string’s probability and its quality. Motivated by this intuition, sampling methods like top-${k}$ (Fan et al., 2018) and nucleus sampling (Holtzman et al., 2020) skew a language model towards high-probability strings. Indeed, these methods dramatically improve the quality of text sampled from the model (Wiher et al., 2022). Over the years, several studies have since contributed to a better understanding of the nuances in the probability–quality relationship (Holtzman et al., 2020; Zhang et al., 2021; Basu et al., 2021; Meister et al., 2022) and led to the development of more sophisticated sampling methods (Basu et al., 2021; Hewitt et al., 2022; Meister et al., 2023b).
@@
This paper seeks to explain the relationship between probability and quality in the specific case of aligned language models, i.e., language models explicitly aligned with human preferences, e.g., through Reinforcement Learning with Human Feedback (RLHF; Leike et al., 2018; Ziegler et al., 2020; Stiennon et al., 2020; Ouyang et al., 2022a, b; Korbak et al., 2022, 2023) In particular, we provide a formal argument and empirical evidence that in RLHF-tuned models, there is an anti-correlation, i.e., a trade-off, between probability and quality. An implication of this finding is that sampling adaptors (Meister et al., 2023a)—post-hoc modifiers of token probabilities at each decoding time step—can control this trade-off by modifying string probability of the generated text.
@@
In the theoretical portion of this paper, we formalize the probability–quality trade-off in aligned language models. Specifically, we show that for corpora of generated strings of a large enough size, the average log-probability under ${p}$ trades off with the average score assigned by a reward model (Gao et al., 2023). We call this a probability–quality trade-off because the reward score is often used to rank text based on various notions of quality (Perez et al., 2022; Lee et al., 2024). This trade-off follows straightforwardly from concentration inequalities and can be seen as a direct corollary of the asymptotic equipartition property (AEP; Cover and Thomas, 2006). In addition to the standard version of the AEP based on Chebyshev’s inequality, we also prove a tighter version of the AEP based on a Chernoff bound that applies for certain language models, including Transformer-based models (Vaswani et al., 2017). Lastly, we show that when used with RLHF-tuned models, sampling adaptors directly influence the probability of a generated string under ${p}$, and in so doing, determine the average quality of generated text by choosing a point on this trade-off. Interestingly, this trade-off predicts the emergence of Simpson’s paradox, which we also observe in our experiments.
@@
In the empirical part of this paper, we present two sets of experiments. First, with synthetic data, we construct toy language and reward models to validate our theoretical results. This gives easily reproducible empirical evidence for our claim. Then, with a second set of experiments, we show that this trade-off exists in practice with open-sourced RLHF-tuned models, and that common sampling adaptors allow us to control where a corpus of generated text will lie on the trade-off.
@@ 2 The Probability–Quality Relationship
Let ${{\Sigma}}$ denote a vocabulary and ${\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}$ denote a string from its Kleene closure, i.e., the set of all finite strings constructed from tokens ${y}\in{{\Sigma}}$. A string’s probability under a language model ${p}({\boldsymbol{y}})$ is commonly used to evaluate its quality with the underlying assumption that a higher probability string should be more human-like, i.e., there is a positive correlation between string probability and quality. [^1] Probability here refers to string likelihood under a general-purpose, unconditional language model. The distinction is important since we will examine the relationship between this probability and the quality of strings in an aligned language model. For example, a considerable number of studies on language modeling methods report measures of perplexity to quantify the quality of text produced by the model (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020). On the other hand, this assumption has been challenged by various works, culminating in what has become the “probability quality paradox” (Zhang et al., 2021; Meister et al., 2022).
The probability quality paradox states that a string’s probability is positively correlated with its quality up to an inflection point, after which it becomes negatively correlated. Meister et al. (2022) show that this inflection point lies near the entropy of a language model trained on human text, and argue that high-quality text has the same information content as natural language strings. To that end, the paradox has inspired various sampling schemes like locally typical (Meister et al., 2023b) and $\eta$ sampling (Hewitt et al., 2022).
In this paper, we investigate the probability–quality relationship in aligned models for two reasons. First, they have an additional constraint—they are fine-tuned to only produce high-quality text—and it is unclear how this might influence the probability–quality relationship. Second, aligned language models share parallels to the conditional language models often found in machine translation and controlled text generation, for which prior work has found relationships not seen in unconditional language models (Callison-Burch et al., 2007; Banchs et al., 2015; Teich et al., 2020; Sulem et al., 2020; Lim et al., 2024).
@@ 3 Learning from Human Feedback
Because our investigation focuses on aligned language models, we now introduce RLHF—a popular alignment method. For a given text generation task, we are interested in producing text that is aligned to human preferences for the task. Let ${\mathcal{A}}={\{{\textsc{--}},{\textsc{+}}\}}$ – , + } denote binary judgments of alignment, and ${A}$ be an ${\mathcal{A}}$-valued random variable. Then, we can formalize the goal of alignment as obtaining an aligned language model ${q_{{\textsc{+}}}}({\boldsymbol{y}})$ close to the true human-aligned distribution over strings ${{p}_{{\textsc{+}}}}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{p}({\boldsymbol{y}}\mid{A}={\textsc{+}})$, such that strings with high probability under ${q_{{\textsc{+}}}}$ also receive positive scores from human annotators. For example, strings that are offensive should have lower probability under ${q_{{\textsc{+}}}}$ for chat-related tasks.
RLHF is a widely used method of finding such a model. At the core of RLHF is a reward function, which models preferences of human annotators. Formally, it can be denoted as ${{r}}\colon{{{{\Sigma}}^{*}}}\rightarrow(-\infty,{{B}}]$ where ${{B}}$ is a bound in $\in{\mathbbm{R}_{+}}$. [^2] We assume that the reward function is bounded, following Levine (2018); Korbak et al. (2022). In practice, the reward function is typically parameterized by a neural network and derived by modeling preferences with a Bradley–Terry model (Bradley and Terry, 1952) and a ranked dataset. The human-aligned language model ${{p}_{{\textsc{+}}}}({\boldsymbol{y}})$, reward function ${{r}}({\boldsymbol{y}})$ and prior language model ${p}({\boldsymbol{y}})$ can be related as follows (Korbak et al., 2022):
${{p}_{{\textsc{+}}}}({\boldsymbol{y}})=\frac{{p}({\boldsymbol{y}})\exp\left(\frac{1}{\beta}{{r}}({\boldsymbol{y}})\right)}{{{Z}({\textsc{+}})}}$ (1)
where $\beta\in{\mathbbm{R}_{+}}$ is a scaling factor and
${{Z}({\textsc{+}})}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{\sum_{{\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}}}{p}({\boldsymbol{y}})\exp\left(\frac{1}{\beta}{{r}}({\boldsymbol{y}})\right)$ (2)
@@
This objective can also be seen as KL-regularized reward learning (Stiennon et al., 2020).
@@
The implication here is that the result we prove for RLHF-tuned models can be trivially extended to any conditionally aligned language model, like the ones often seen in controlled text generation (Hu et al., 2017; Krause et al., 2021; Yang and Klein, 2021; Liu et al., 2021; Zhang et al., 2023).
@@ 4 Sampling Adaptors
Text generation is often performed by sampling from a probabilistic language model ${p}({\boldsymbol{y}})$, also commonly referred to as decoding. Sampling is usually performed autoregressively, where a token ${\overline{{y}}}\in{\overline{{{\Sigma}}}}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{{\Sigma}}\cup{\{{\textsc{eos}}\}}$ is iteratively sampled from ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$ at each time step $t$ until the special end-of-sequence eos token is reached. Sampling adaptors (Meister et al., 2022) are post-hoc alterations of ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$ that have been shown to dramatically improve the quality of text produced by language models, and are often considered an integral part of a text generation pipeline (Wiher et al., 2022). Common examples of sampling adaptors include top-${k}$ (Fan et al., 2018) and nucleus sampling (Holtzman et al., 2020).
Formally, we can define a sampling adaptor ${\boldsymbol{\alpha}}\colon{{\Delta}^{\mid{\overline{{{\Sigma}}}}\mid-1}}\rightarrow{{\mathbbm{R}_{+}}^{\mid{\overline{{{\Sigma}}}}\mid-1}}$ as a function from a distribution over ${\overline{{{\Sigma}}}}$ to an unnormalized distribution. Sampling adaptors are applied pointwise, such that
${\tilde{{p}}}(\cdot\mid{{\boldsymbol{y}}_{<t}})\propto{\boldsymbol{\alpha}}({p}(\cdot\mid{{\boldsymbol{y}}_{<t}}))$ (5)
where ${\tilde{{p}}}$ denotes the resultant distribution over ${\overline{{{\Sigma}}}}$. The probability of a string when a sampling adaptor ${\boldsymbol{\alpha}}$ is applied to a language model ${p}$ is then:
@@
${\tilde{{p}}}({\boldsymbol{y}})=\frac{{\boldsymbol{\alpha}}\big{(}{p}(\cdot\mid{\boldsymbol{y}})\big{)}({\textsc{eos}})\prod^{|{\boldsymbol{y}}|}_{t=1}{\boldsymbol{\alpha}}\big{(}{p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\big{)}({{y}_{t}})}{{{Z}_{\boldsymbol{\alpha}}({\boldsymbol{y}})}}$ (6)
where ${{Z}_{\boldsymbol{\alpha}}({\boldsymbol{y}})}$ is the normalizing constant. [^3] We note that this description of sampling adaptors is slightly different to Meister et al.’s (2023a), where ${\boldsymbol{\alpha}}$ is instead defined as a function that produces a normalized distribution over ${\overline{{{\Sigma}}}}$. We make this choice to simplify our exposition in later sections because the two descriptions are equivalent.
In general, the application of a sampling adaptor ${\boldsymbol{\alpha}}$ to a language model ${p}$ induces a different distribution, i.e., ${\tilde{{p}}}\neq{p}$. For example, top-${k}$ sampling is designed to produce text that has higher probability under ${p}({\boldsymbol{y}})$, and for ${k}<|{\overline{{{\Sigma}}}}|$ it is easy to see that ${\tilde{{p}}}\neq{p}$. As it turns out, ensuring that the resultant distribution converges to the underlying model will become important in § 5.2.
@@
To correct the distribution, we can use the Independent Metropolis-Hastings Algorithm (IMHA; Metropolis et al., 1953; Hastings, 1970; Wang, 2022). The IMHA is a Markov Chain Monte Carlo (MCMC) method that simulates sampling from a target distribution using a proposal distribution. The idea is that by sampling sequentially from the proposal, i.e., ${\tilde{{p}}}$ and appropriately accepting or rejecting samples, the generated Markov chain (i.e., the sequence of samples) converges to a stationary distribution equal to the target distribution, i.e., ${p}$. Convergence is achieved when there is no autocorrelation between consecutive samples. In the case of categorical variables like strings, this can be measured with Cramer’s V (Ialongo, 2016; Deonovic and Smith, 2017). We refer the reader to Wang (2022) for a formal treatment and detail the IMHA’s acceptance–rejection protocol in Algorithm 1.
Algorithm 1 The IMH Algorithm
Corpus size $N$, initial sample ${\boldsymbol{y}}\sim{\tilde{{p}}}$, language model ${p}$, model with sampling adaptor ${\tilde{{p}}}$
Sampled corpus ${\mathcal{Y}}$ of size $N$
${\mathcal{Y}}\leftarrow\{\}$
for $step=1$ to $N$ do
     ${{\boldsymbol{y}}^{\prime}}\sim{\tilde{{p}}}$
     ${a}({\boldsymbol{y}},{{\boldsymbol{y}}^{\prime}})\leftarrow\min\left(\frac{{p}({{\boldsymbol{y}}^{\prime}})\,{\tilde{{p}}}({\boldsymbol{y}})}{{p}({\boldsymbol{y}})\,{\tilde{{p}}}({{\boldsymbol{y}}^{\prime}})},1\right)$
     $r\sim U(0,1)$
     if ${a}>r$ then
         ${{\boldsymbol{y}}^{\prime}};{\mathcal{Y}}\leftarrow{\mathcal{Y}}\cup\{{{\boldsymbol{y}}^{\prime}}\}$
         ${\boldsymbol{y}}\leftarrow{\boldsymbol{y}}^{\prime}$
     else
         ${{\boldsymbol{y}}^{\prime}};{\mathcal{Y}}\leftarrow{\mathcal{Y}}\cup\{{\boldsymbol{y}}\}$
     end if
end for
Understanding the effectiveness and role of sampling adaptors in a text generation pipeline is an active area of research. Early sampling adaptors were often found through trial and error and motivated by intuitive explanations rather than formal arguments. The success of top-${k}$ and nucleus sampling, for example, is often informally explained by how they avoid the poorly modeled long tail by reallocating that probability mass to better-modeled tokens. More recently, Basu et al. (2021) and others (Holtzman et al., 2020; Hewitt et al., 2022; Meister et al., 2023b) have begun to treat sampling adaptors more formally.
In this paper, we build on the work of Meister et al. (2023a), who analyze sampling adaptors through the lens of a precision–recall trade-off. Specifically, they argue that sampling adaptors realign the prior ${p}({\boldsymbol{y}})$ to one that emphasizes precision [^4] Precision refers to the generalization of the term to generative modeling. See Sajjadi et al. (2018); Djolonga et al. (2020). with respect to the underlying data-generating distribution, thus resulting in higher quality text being produced on average. Though we find a different trade-off in this work, in § 5.2 we similarly show how sampling adaptors can control the average reward of text sampled from a language model.

### v2

@@ Abstract
The relationship between the quality of a string, as judged by a human reader, and its probability, ${p}({\boldsymbol{y}})$ under a language model undergirds the development of better language models. For example, many popular algorithms for sampling from a language model have been conceived with the goal of manipulating ${p}({\boldsymbol{y}})$ to place higher probability on strings that humans deem of high quality (Fan et al., 2018; Holtzman et al., 2020). In this article, we examine the probability–quality relationship in language models explicitly aligned to human preferences, e.g., through reinforcement learning through human feedback. We show that, when sampling corpora from an aligned language model, there exists a trade-off between the strings’ average reward and average log-likelihood under the prior language model, i.e., the same model before alignment with human preferences. We provide a formal treatment of this phenomenon and demonstrate how a choice of sampling adaptor allows for a selection of how much likelihood we exchange for the reward.
@@ 1 Introduction
The relationship between the probability of a string and its quality as judged by a human reader is fundamental to the goal of producing high-quality text from probabilistic language models. For example, the belief that a string ${\boldsymbol{y}}$ with a higher probability under a language model ${p}$ should be of higher quality (Graves, 2012; Fan et al., 2018; Holtzman et al., 2020; Zhang et al., 2021) has motivated the development of sampling adaptors like top-${k}$ (Fan et al., 2018) and nucleus sampling (Holtzman et al., 2020). The intuition behind this belief is that, under a well-calibrated language model trained primarily on human-written text, strings that occur with high probability should be more human-like, and thus judged by humans to be of higher quality. Indeed, sampling methods that skew a language model towards high-probability strings have been shown to dramatically improve the quality of text sampled from a model (Wiher et al., 2022). Over the years, several studies have since contributed to a more nuanced understanding of the probability–quality relationship (Holtzman et al., 2020; Zhang et al., 2021; Basu et al., 2021) and led to the development of more sophisticated sampling methods (Hewitt et al., 2022; Meister et al., 2022).
@@
This paper seeks to explain the relationship between probability and quality in the specific case of aligned language models, i.e., language models explicitly fine-tuned to better match human preferences, e.g., through reinforcement learning with human feedback (RLHF; Christiano et al., 2017; Leike et al., 2018; Ziegler et al., 2020; Stiennon et al., 2020; Ouyang et al., 2022a, b; Korbak et al., 2022, 2023). In particular, we provide a formal argument and empirical evidence that in RLHF-tuned models, there is an inherent anti-correlation, i.e., a trade-off, between probability and quality. We introduce a new class of sampling methods with useful theoretical properties, globally normalized sampling adaptors, and show that they can be used to control this trade-off by modifying string probability of the generated text.
@@
In the theoretical portion of this paper, we formalize the probability–quality trade-off in aligned language models. Specifically, we show that for corpora of generated strings of a large enough size, the average log-probability under the prior (unaligned) language model ${p}$ trades off with the average score assigned by a reward model (Gao et al., 2023). We refer to this phenomenon as a probability–quality trade-off because, in the context of aligned language models, quality is often operationalized by the reward model (Perez et al., 2022; Lee et al., 2024). This trade-off follows straightforwardly from an application of a concentration inequality and can be seen as a direct corollary of the asymptotic equipartition property (AEP; Cover and Thomas, 2006). In addition to the standard version of the AEP based on Chebyshev’s inequality, we also prove a tighter version of the AEP based on a Chernoff bound that applies to certain language models, including Transformer-based models (Vaswani et al., 2017). Lastly, we show how an appropriate choice of globally normalized sampling adaptor can control the probability of a generated string under the prior ${p}$. The implication of this finding is that they can be used to choose a point on the probability–quality trade-off and control the average quality of generated text. Interestingly, the trade-off predicts the emergence of Simpson’s paradox, which we also observe in our experiments.
@@
In the empirical part of our paper, we present two sets of experiments. First, with synthetic data, we construct toy language and reward models to further demonstrate our theoretical results. This gives easily reproducible empirical evidence for our claim. Then, with a second set of experiments, we show that this trade-off exists in practice with open-sourced RLHF-tuned models, and that globally normalized sampling adaptors do allow us to control where a corpus of generated text will lie on the trade-off. Notably, we find that in practice, locally normalized sampling adaptors (e.g., nucleus sampling; Holtzman et al., 2020; Meister et al., 2023a) are sufficiently close approximations of their globally normalized counterparts, and can be effectively used to control the trade-off.
@@ 2 The Probability–Quality Relationship
Let ${{\Sigma}}$ be an alphabet. A string ${\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}$ is a finite list of symbols drawn from ${{\Sigma}}$. The set of all strings is called ${{\Sigma}}$’s Kleene closure and is denoted as ${{{{\Sigma}}^{*}}}$. A language model ${p}$ is a probability distribution over ${{{{\Sigma}}^{*}}}$. It is often assumed that there is a positive correlation between the quality of a string and its probability under a language model trained on human text ${p}({\boldsymbol{y}})$ (Graves, 2012; Fan et al., 2018; Holtzman et al., 2020; Zhang et al., 2021). [^1] Probability here refers to string likelihood under a general-purpose, unconditional language model. The distinction is important since we will examine the relationship between this probability and the quality of strings in an aligned language model. The idea behind this assumption is that under such a model, a higher probability string should be more human-like, and thus judged by humans to be of higher quality. Because of this positive correlation, string probability is commonly used to reason about text quality in the context of natural language processing. For example, a considerable number of studies on language modeling methods report measures of perplexity to quantify the quality of text produced by the model (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020; Pillutla et al., 2023).
More recently, several authors have observed evidence that this positive correlation breaks down at extremes, finding instead that a string’s probability is positively correlated with its quality up to an inflection point, after which it becomes negatively correlated (Yang et al., 2018; Stahlberg and Byrne, 2019; Holtzman et al., 2020; Zhang et al., 2021; Meister et al., 2022). Meister et al. (2022) show that this inflection point lies near the entropy of human-written text, and hypothesize that a string should encode a similar amount of information to natural language strings to be considered high-quality. This finding has since inspired various sampling adaptors, e.g., locally typical (Meister et al., 2023b) and $\eta$ sampling (Hewitt et al., 2022).
In this paper, we investigate the probability–quality relationship in aligned language models. There are two reasons for this choice. First, aligned models have an additional constraint—they are fine-tuned to only produce high-quality text (as judged by a reward model). And, a priori, it is unclear how alignment might influence the probability–quality relationship. Second, aligned language models are mathematically similar to the conditional language models often found in machine translation and controlled text generation, for which prior work has found relationships not seen in unconditional language models (Callison-Burch et al., 2007; Banchs et al., 2015; Teich et al., 2020; Sulem et al., 2020; Lim et al., 2024).
@@ 3 Learning from Human Feedback
Because our investigation focuses on aligned language models, we now introduce RLHF, a popular algorithm for alignment. Given a task of interest, we are interested in producing strings that are aligned with human preferences for that task. Let ${\mathcal{A}}={\{{\textsc{--}},{\textsc{+}}\}}$ – , + } denote binary judgments of alignment, and ${A}$ be an ${\mathcal{A}}$-valued random variable. Then, we assume the existence of a true human-aligned distribution over strings ${{p}_{{\textsc{+}}}}({\boldsymbol{y}})\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{p}({\boldsymbol{y}}\mid{A}={\textsc{+}})$, such that strings with high probability under ${{p}_{{\textsc{+}}}}$ also receive positive scores from human annotators. For example, we may desire that strings that are offensive should have lower probability under ${{p}_{{\textsc{+}}}}$ for chat-related tasks. With this, we can formalize the goal of alignment as obtaining an aligned language model ${q_{{\textsc{+}}}}({\boldsymbol{y}})$ such that ${q_{{\textsc{+}}}}({\boldsymbol{y}})$ is a good approximation to ${{p}_{{\textsc{+}}}}({\boldsymbol{y}})$.
RLHF is a widely used method of finding such a model. At the core of RLHF is a reward function, which models preferences of human annotators, denoted as ${{r}}\colon{{{{\Sigma}}^{*}}}\rightarrow(-\infty,{{B}}]$ where ${{B}}$ is a bound in $\in{\mathbbm{R}_{>0}}$. [^2] We assume that the reward function is bounded, following Levine (2018); Korbak et al. (2022). In practice, the reward function is itself typically parameterized by a neural network and derived by modeling preferences with a Bradley–Terry model (Bradley and Terry, 1952) and a ranked dataset. The human-aligned language model ${{p}_{{\textsc{+}}}}({\boldsymbol{y}})$, reward function ${{r}}({\boldsymbol{y}})$ and prior language model ${p}({\boldsymbol{y}})$ can be related as follows (Korbak et al., 2022):
${{p}_{{\textsc{+}}}}({\boldsymbol{y}})=\frac{{p}({\boldsymbol{y}})\exp\left(\frac{1}{\beta}{{r}}({\boldsymbol{y}})\right)}{{{Z}({\textsc{+}})}},$ (1)
where $\beta\in{\mathbbm{R}_{>0}}$ is a scaling factor and
${{Z}({\textsc{+}})}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{\sum_{{\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}}}{p}({\boldsymbol{y}})\exp\left(\frac{1}{\beta}{{r}}({\boldsymbol{y}})\right).$ (2)
@@
This objective can be seen as KL-regularized reinforcement learning (Stiennon et al., 2020).
@@
The implication of this is that the result we prove for RLHF-tuned models can be trivially extended to any conditionally aligned language model, like the ones often deployed in controlled text generation (Hu et al., 2017; Krause et al., 2021; Yang and Klein, 2021; Liu et al., 2021; Zhang et al., 2023).
@@ 4 Sampling Adaptors
We are often interested in sampling strings from a language model ${p}({\boldsymbol{y}})$. Sampling is usually performed autoregressively, i.e., we sample a token ${\overline{{y}}}\in{\overline{{{\Sigma}}}}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{{\Sigma}}\cup{\{{\textsc{eos}}\}}$ iteratively from ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$ at each time step $t$ until the special end-of-sequence eos token is reached. Locally normalized sampling adaptors (Meister et al., 2023a) are post-hoc alterations of ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$ that have been shown to dramatically improve the quality of text produced by language models, and are often considered an integral part of a text generation pipeline (Wiher et al., 2022). Common examples include top-${k}$ (Fan et al., 2018) and nucleus sampling (Holtzman et al., 2020).
Applying a locally normalized sampling adaptor to ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$ at each $t$ is commonly formulated as the composition of two steps. First, the application of a transform function ${\boldsymbol{\gamma}}\colon{{\Delta}^{\mid{\overline{{{\Sigma}}}}\mid-1}}\rightarrow{{\mathbbm{R}_{>0}}^{\mid{\overline{{{\Sigma}}}}\mid}}$ that maps the distribution over ${\overline{{{\Sigma}}}}$ to a vector of non-negative values. The transform function is responsible for the core logic of the sampling adaptor, e.g., assigning tokens outside the top-${k}$ zero probability. Second, a normalization step is performed to ensure a valid distribution over ${\overline{{{\Sigma}}}}$, i.e, one where ${\sum_{{\overline{{y}}}\in{\overline{{{\Sigma}}}}}}{\boldsymbol{\gamma}}({p}({\overline{{y}}}\mid{{\boldsymbol{y}}_{<t}}))=1$. When a locally normalized sampling adaptor is applied to a language model ${p}$, we can express the induced distribution ${\ddot{{p}}}$ as
${\ddot{{p}}}({\boldsymbol{y}})$
$=\frac{{\boldsymbol{\gamma}}({p}({\textsc{eos}}\mid{\boldsymbol{y}}))}{{\sum_{{\overline{{y}}}\in{\overline{{{\Sigma}}}}}}{{\boldsymbol{\gamma}}({p}({\overline{{y}}}\mid{\boldsymbol{y}}))}}\frac{{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\boldsymbol{\gamma}}({p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}}))}{{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\sum_{{\overline{{y}}}\in{\overline{{{\Sigma}}}}}}{\boldsymbol{\gamma}}({p}({\overline{{y}}}\mid{{\boldsymbol{y}}_{<t}}))}$ (5)
Despite their empirical success, locally normalized sampling adaptors have several undesirable theoretical properties. First, they can induce a language model that is not tight, i.e., one where probability mass can be placed on infinite-length strings (Welleck et al., 2020; Du et al., 2023). For instance, this can occur if zero probability is placed on eos at every step $t$. Second, the normalization operation at each $t$ can produce unexpected behavior. Specifically, because the denominator used to normalize the local distribution at each $t$ is dependent on the probability mass assigned to other tokens, a string assigned a higher score with the transform function is not necessarily assigned a higher probability under the induced language model ${\ddot{{p}}}$. This can lead to less desirable strings being sampled more frequently. See App. A for an example.
We thus introduce globally normalized sampling adaptors, a procedure to sample from a language model that circumvents these issues. Globally normalized sampling adaptors are also defined with respect to a transform function ${\boldsymbol{\gamma}}$, and thus every existing locally normalized sampling adaptor can be mapped to its globally normalized counterpart. The transform function is similarly applied pointwise at each time step to ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$. But without the normalization step, we define the probability of a string under the induced model ${\widetilde{{p}}}$ as
@@
${\widetilde{{p}}}({\boldsymbol{y}})=\frac{{\boldsymbol{\gamma}}\big{(}{p}(\cdot\mid{\boldsymbol{y}})\big{)}({\textsc{eos}}){\prod^{|{\boldsymbol{y}}|}_{t=1}}{\boldsymbol{\gamma}}\big{(}{p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\big{)}({{y}_{t}})}{{{Z}_{\boldsymbol{\gamma}}({{\Sigma}})}}.$ (6)
where the normalizing constant ${{Z}_{\boldsymbol{\gamma}}({{\Sigma}})}$ is defined
${{Z}_{\boldsymbol{\gamma}}({{\Sigma}})}$
$={\sum_{{\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}}}{\boldsymbol{\gamma}}\big{(}{p}(\cdot\mid{\boldsymbol{y}})\big{)}({\textsc{eos}}){\prod^{|{\boldsymbol{y}}|}_{t=1}}{\boldsymbol{\gamma}}\big{(}{p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\big{)}({{y}_{t}}).$ (7)
The language model induced by a globally normalized sampling adaptor is always tight. Under the assumption that ${{Z}_{\boldsymbol{\gamma}}({{\Sigma}})}$ is finite (Cotterell et al., 2023), it is easy to see that ${\sum_{{\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}}}{\widetilde{{p}}}({\boldsymbol{y}})=1$. It is also easy to see that
${\widetilde{{p}}}({\boldsymbol{y}})\propto{\boldsymbol{\gamma}}\big{(}{p}(\cdot\mid{\boldsymbol{y}})\big{)}({\textsc{eos}}){\prod^{|{\boldsymbol{y}}|}_{t=1}}{\boldsymbol{\gamma}}\big{(}{p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\big{)}({{y}_{t}}).$ (8)
That is, the probability of a string under the adapted model is determined only by the transform function, which makes it convenient to reason about how a choice of transform function can modify the properties of the induced distribution ${\widetilde{{p}}}$. Importantly, we use this property in § 5.2 to show that globally normalized sampling adaptors can be used to control the probability–quality trade-off.
@@
To generate strings from a language model using a globally normalized sampling adaptor, we use the Independent Metropolis–Hastings Algorithm (IMHA; Metropolis et al., 1953; Hastings, 1970; Wang, 2022).333 Directly sampling from ${\widetilde{{p}}}$ is not possible because it requires computing the constant ${{Z}_{\boldsymbol{\gamma}}({{\Sigma}})}$. This sum over the countably infinite set of finite strings is in practice unknown. Using IMHA avoids this issue as it only requires the knowledge of the target distribution up to a multiplicative constant. The IMHA is a Markov Chain Monte Carlo (MCMC) method that simulates sampling from a target distribution using a proposal distribution. We detail how the IMHA’s acceptance–rejection protocol can be used to generate text with a globally normalized sampling adaptor in Algorithm 1. The idea is that by sampling sequentially from the proposal (i.e., the distribution induced by a locally normalized sampling adaptor, ${\ddot{{p}}}$) and appropriately accepting or rejecting samples, the generated Markov chain (i.e., the sequence of samples) converges to a stationary distribution equal to the target distribution (i.e., the one induced by a globally normalized sampling adaptor, ${\widetilde{{p}}}$). Convergence is often diagnosed by the absence of autocorrelation between consecutive samples (Deonovic and Smith, 2017). In the case of categorical variables like strings, this can be measured with Cramér’s V (Ialongo, 2016).
1.def IMHA $\left(N,{\widetilde{{p}}},{\ddot{{p}}}\right)$:
2.   ${\boldsymbol{y}}\sim{\ddot{{p}}}$
3.   ${\mathcal{Y}}\leftarrow\{\}$
4.   for $n\leftarrow 1\ldots N$ :
5.      ${{\boldsymbol{y}}^{\prime}}\sim{\ddot{{p}}}$
6.      ${a}({\boldsymbol{y}},{{\boldsymbol{y}}^{\prime}})\leftarrow\frac{{\widetilde{{p}}}({{\boldsymbol{y}}^{\prime}})\,{\ddot{{p}}}({\boldsymbol{y}})}{{\widetilde{{p}}}({\boldsymbol{y}})\,{\ddot{{p}}}({{\boldsymbol{y}}^{\prime}})}$
7.      $r\sim U(0,1)$
8.      if ${a}>r$ :
9.         ${\mathcal{Y}}\leftarrow{\mathcal{Y}}\cup\{{{\boldsymbol{y}}^{\prime}}\}$
10.         ${\boldsymbol{y}}\leftarrow{\boldsymbol{y}}^{\prime}$
11.      else
12.         ${\mathcal{Y}}\leftarrow{\mathcal{Y}}\cup\{{\boldsymbol{y}}\}$
13.   return ${\mathcal{Y}}$
Algorithm 1 IMH Algorithm.